OpenStudy (anonymous):

What's the reminder when 2^34 is divided by 5? plz help

OpenStudy (isaiah.feynman):

2^34 is a very large number, when you divide it by 5 you have no remainder.

OpenStudy (akashdeepdeb):

I am not sure of any direct way to do this but yeah there is a logical way! We OBVIOUSLY will not calculate the 2^34 value :P So what we do is we find out what are the remainders at each stage and find a pattern! Like this.... 2^1 leaves remainder --- 2 --- when divided by 5 2^2 leaves remainder --- 4----when divided by 5 2^3 leaves remainder --- 3--- when divided by 5 2^4 leaves remainder --- 1----when divided by 5 2^5 leaves remainder --- 2 --- when divided by 5 2^6 leaves remainder --- 4----when divided by 5 And it goes on so we see that! The pattern of remainder is 2,4,3,1,2,4,3,1..... All we have to do is see which one corresponds to 34!! :) Understood? :)

OpenStudy (anonymous):

31

OpenStudy (zzr0ck3r):

so 2 has order 4 mod(5)

OpenStudy (akashdeepdeb):

@Isaiah.Feynman Try it out with a calc! It does have a remainder! :)

OpenStudy (anonymous):

4 i think

OpenStudy (isaiah.feynman):

Looks like my its beyond my calculator. lol

OpenStudy (akashdeepdeb):

XD

OpenStudy (anonymous):

3+1

OpenStudy (zzr0ck3r):

2^{34} conruent 2^6 congruent to 4 mod(5)

OpenStudy (akashdeepdeb):

Exactly so it should be! 4 !! :D Because 2^36 remainder is 1 [Because 36 is a multiple of 4] And so 2^35 would be 3 And so 2^34 would be 4!! Check the pattern @CoolJolie

OpenStudy (anonymous):

how ur saying 2^36 is four@AkashdeepDeb

OpenStudy (anonymous):

sorry 1

OpenStudy (akashdeepdeb):

Because 2^36 has a degree of 36 and we know that 36 is a multiple of 4! And we saw that 2^4 had a remainder of 1 too!! :) And so did 2^8 = 256 has a remainder of 1 :D Got it? :) And so we trace the pattern backwards and we get 2^34 as 4 :D

OpenStudy (anonymous):

yes ur right

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

i thought 2^3=3 2^4=1 so 3+1=4 and

OpenStudy (akashdeepdeb):

Haha. No! That is a funny way though! : ' )

OpenStudy (anonymous):

but 2^36= 3+4=7 2^7=1

OpenStudy (anonymous):

bcoz 2^4 leaves 1 when divided by 5 hence 2^32=(2^4)^8 will leave remainder 1 when divided by 5 hence 2^34=2^2 * 2^32 so when 2^34 is divided by 5 remainder is 4